# Fair Uncertainty Quantification for Depression Prediction

### v1

@@ I Introduction
As research delves deeper, despite the outstanding capabilities for learning representations, researchers have found that the deep learning models are untrustworthy. One major reason is that deep models are typically ”black boxes” and lack interpretability. Additionally, observations have shown that predictions often exhibit ”overconfidence.” In depression prediction, [22, 11] address this issue through uncertainty quantification (UQ). Calibrated predictions with UQ not only provide accurate predictions but also output predictions with the confidence level.
Moreover, few studies have focused on the issue that data-driven deep learning algorithms are always prone to predictive bias, which varies across different groups and leads to unfair predictions. Such biases can come from various sources, including imbalanced data distribution, inherent algorithmic design, and task-specific requirements [23, 24, 25]. Empirical evidence shows that depression prediction is unfair across demographic groups, and such bias directly weakens the validity of uncertainty quantification, as shown in Fig. 1(a). [26, 27] seek to achieve fair depression prediction by eliminating biases associated with specific attributes, but they overlook uncertainty quantification and calibration. However, achieving fair depression prediction within a UQ framework is nontrivial. First, incorporating fairness considerations into existing UQ methods for depression recognition is neither intuitive nor straightforward. Most prevailing approaches prioritize aggregate statistical validity and, in doing so, neglect the demographic stratifications with sensitive attributes. Consequently, their uncertainty estimates tend to privilege majority groups while failing to adequately cover minority groups, thereby introducing unfairness. Second, stratifying the data by sensitive attributes substantially reduces the sample size within each group, which can, in turn, impair the statistical validity of the resulting UQ estimates. In this paper, we consider reliable depression prediction through effective UQ and fairness-aware optimization. The motivation for this work is summarized as follows:
@@ III-B UQ for depression prediction
Referring to [11], we adopt Conformalized Quantile Regression (CQR) [45] for uncertainty quantification in depression prediction. Specifically, during the training, we apply the pinball loss function as the learning objective to achieve quantile prediction for depression scores.
@@ IV-C Results
In Table I, we evaluate several commonly used depression recognition algorithms on widely adopted depression datasets. Specifically, for the AVEC 2013 & 2014 datasets, we focus on the video modality, whereas for DAIC-WOZ and MODMA, we use the audio modality for predictions. Among these datasets, AVEC and DAIC-WOZ contain only gender as a sensitive attribute, while MODMA includes both gender and age. Due to the number of participants in MODMA and the more evident prediction biases observed, we categorize age into two groups: ’Young’ representing participants under 30 years old, and ’Old’ representing those over 30. Experimental results indicate that biases exist in depression prediction concerning both age and gender. However, the manifestation of these biases varies across datasets. Specifically, in the video modality, there is no consistent pattern of higher or lower prediction errors across genders. Given the limitations of small datasets and the lack of pathological support for depression, we can only conclude that prediction biases are present in depression recognition. However, there is currently insufficient justification or evidence to warrant bias elimination.
In Tables II and III, we conduct fair depression prediction for two sensitive attributes: gender and age. This is achieved by performing interval prediction using both the overall calibration set and groups with different sensitive attributes (s = 0 or 1). However, partitioning the calibration set reduces the sample size available for calibration, which in turn affects the validity of CP. This limitation is evident in CP for the AVEC 2013 dataset, as well as in CP, CQR, and CFQR for the AVEC 2014 dataset. While BFQR maintains overall validity and achieves a certain level of EOC fairness, its optimization strategy prioritizes overall validity rather than fairness, leading to suboptimal fairness performance. In contrast, our proposed method preserves overall validity while explicitly enforcing EOC fairness. Across different datasets, interval prediction methods, and sensitive attributes, our approach consistently achieves optimal fairness in prediction.

### v2

@@ I Introduction
As research delves deeper, despite the outstanding capabilities for learning representations, researchers have found that the deep learning models are untrustworthy. One major reason is that deep models are typically ”black boxes” and lack interpretability. Additionally, observations have shown that predictions often exhibit ”overconfidence”. In depression prediction, [22, 11] address this issue through uncertainty quantification (UQ). Calibrated predictions with UQ not only provide accurate predictions but also output predictions with the confidence level.
Moreover, few studies have focused on the issue that data-driven deep learning algorithms are always prone to predictive bias, which varies across different groups and leads to unfair predictions. Such biases can come from various sources, including imbalanced data distribution, inherent algorithmic design, and task-specific requirements [23, 24, 25]. Kuzucu et al. [26] show that models produce systematically biased uncertainty estimates across demographic groups. Furthermore, empirical evidence shows that depression prediction is unfair across demographic groups, and such bias directly weakens the validity of uncertainty quantification, as shown in Fig. 1(a). [27, 28] seeks to achieve fair depression prediction by eliminating biases associated with specific attributes, but they overlook uncertainty quantification and calibration. However, achieving fair depression prediction within a UQ framework is nontrivial [26]. First, incorporating fairness considerations into existing UQ methods for depression recognition is neither intuitive nor straightforward. Most prevailing approaches prioritize aggregate statistical validity and, in doing so, neglect the demographic stratifications with sensitive attributes. Consequently, their uncertainty estimates tend to privilege majority groups while failing to adequately cover minority groups, thereby introducing unfairness. Second, stratifying the data by sensitive attributes substantially reduces the sample size within each group, which can, in turn, impair the statistical validity of the resulting UQ estimates. In this paper, we consider reliable depression prediction through effective UQ and fairness-aware optimization. The motivation for this work is summarized as follows:
@@ III-B UQ for depression prediction
Referring to [11], we adopt Conformalized Quantile Regression (CQR) [46] for uncertainty quantification in depression prediction. Specifically, during training, we apply the pinball loss function as the learning objective to achieve quantile prediction for depression scores.
@@ IV-B3 Metrics
(a) AVEC 2013
(b) AVEC 2014
(c) DAIC-WOZ
(d) MODMA
Figure 3: Performance of different fairness methods on gender group across various datasets. At the 90 % confidence level ($\alpha=0.1$), our method yields the smallest group-wise coverage gap.
(a) Gender Group
(b) Age Group
Figure 4: Performance of the various fairness methods on the MOMDA test set across different demographic groups: (a) gender and (b) age. At the 90 % confidence level ($\alpha=0.1$), our method yields the smallest group-wise coverage gap.
@@ IV-C Results
In Table I, we evaluate several commonly used depression recognition algorithms on widely adopted depression datasets. Specifically, for the AVEC 2013 & 2014 datasets, we focus on the video modality, whereas for DAIC-WOZ and MODMA, we use the audio modality for predictions. Among these datasets, AVEC and DAIC-WOZ contain only gender as a sensitive attribute, while MODMA includes both gender and age. Due to the number of participants in MODMA and the more evident prediction biases observed, we categorize age into two groups: ’Young’ representing participants under 30 years old, and ’Old’ representing those over 30. Experimental results indicate that biases exist in heterogeneous biases across both age and gender. In the video modality, for example, neither males nor females consistently incur higher (or lower) prediction errors, and the same absence of a clear pattern is observed for speech. Given the limitations of small datasets and the lack of pathological support for depression, we can only conclude that prediction biases are present in depression recognition; we cannot yet determine whether one gender or age group truly has a higher prevalence of depression. Even so, these prediction biases inevitably undermine the fairness of UQ: groups with smaller errors are over-covered, whereas groups with larger errors are under-covered, leading to unequal validity across demographic subgroups.
Tables II and III offer a systematic comparison of several fairness-aware uncertainty-quantification methods across multiple depression datasets, considering two sensitive attributes: gender (present in every dataset) and age (available only in MOMDA). Although all methods meet the global validity requirement at the 90 % confidence level ($\alpha=0.1$; PICP $\geq$ 90 %), marked coverage disparities appear once performance is disaggregated by demographic group. Classical CP and CQR, which were not designed with fairness constraints, consistently under-cover groups with larger prediction errors and over-cover those with smaller errors, producing the widest validity gap. CFQR attempts to mitigate the biases by adding a demographic-parity (DP) adjustment to CQR. DP fairness requires that the post-calibration predictive distributions of different subgroups be similar, but satisfying this constraint widens the prediction intervals, yields only marginal, statistically insignificant gains, and even introduces excess risk that can increase certain gaps. For instance, although CFQR narrows the validity gap of CQR on the MOMDA (age) and DAIC-WOZ datasets, the same adjustment exacerbates disparities on the remaining datasets because the larger intervals overcompensate for random noise rather than systematic bias. In other words, DP fairness definition designed to mitigate prediction bias is ill-suited for uncertainty fairness. BFQR imposes fairness regularization on UQ, but it prioritizes overall coverage so heavily that the resulting intervals become overly conservative; both PICP and MPIW rise, undermining practicality without truly redressing between-group imbalance. This situation is evident in DAIC-WOZ and MOMDA (gender), where BFQR sacrifices tightness while still failing to equalize validity across attributes (as shown in Fig. 3).
In contrast, our proposed method strikes a more elegant balance: it maintains PICP at the confident level for every dataset, significantly narrows MPIW, and minimizes the coverage gap between male and female or younger and older subgroups, thereby ensuring EOC fairness while preserving tight intervals. Across different datasets, methods, and sensitive attributes, our approach consistently achieves optimal fairness in prediction. Although most publicly available depression datasets include only gender labels, our approach shows the same validity consistency and minimal coverage disparity for both gender and age on the MOMDA dataset (as shown in Fig. 4), underscoring its robustness to multiple sensitive factors. This robustness is crucial for reducing the risk of unfair outcomes in clinical practice.